Kim has invested a certain amount of money in stocks and bonds. The annual return from these investments is $650. If the stocks produce $130 more in returns than the bonds, how much money does Kim receive annually from each type of investment?
Answer by macston(5194)

.
B=return from bonds; S=return from stocks=B+$130
.
B+S=$650
B+B+$130=$650
2B=$520
B=$260
ANSWER 1: Return from bonds is $260.
.
S=B+$130
S=$260+$130
S=$390
ANSWER 2: Return from stocks is $390.
.
CHECK:
B+S=$650
$260+$390=$650
$650=$650
.